

********************************************************************
UARS WEEKLY STATUS REPORT - 17 AUG 2001
********************************************************************
***********
SOLSTICE
***********
With the exception of a brief off period last night (01229) due to TMON
response, SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods and stellar calibration observations
whenever possible. We continue to gather enough data to meet all the
SOLSTICE science objectives during this period of reduced tape recorder
operations.
***********
SUSIM
***********
SUSIM was turned off on Aug. 8 and back on Aug. 11. On Aug. 17, SUSIM
was unexpectedly turned off as were other instruments due to the TMON.
Otherwise, SUSIM has completed its daily solar scans this past week
during periods of predicted UARS telemetry.
With the impending end of UARS operations, we display a summary of
downloads at SUSIM's ftp site.
---
SUMMARY OF SUSIM DATA DOWNLOADS
8/11/01
Since April 2000, there have been a total 2666 downloads of SUSIM data
from 258 unique downloaders. These DO NOT include those just visiting
the SUSIM WWW site, but rather, those actually downloading SUSIM data
from SUSIM's ftp site. Many individuals make repeat visits to get
CURRENT DATA. The downloaders include individuals from the following
institutions and countries.
NASA CENTERS (4)
Ames Research Center
Goddard Space Flight Center
Jet Propulsion Laboratory
Marshall Space Flight Center
US UNIVERSITIES and EDUCATIONAL INSTITUTIONS (20)
Applied Physics Laboratory, Johns Hopkins University
California State University, Northridge
Florida Institute of Technology
New Mexico Institute of Mining and Technology
Rochester Institute of Technology
Southwestern University
State of Indiana, Department of Education
State University of New York, Stony Brook
University at Albany, State University of New York
University of Arizona
University of California, Berkeley
University of Colorado
University of Florida
University of Maryland
University of Michigan
University of North Dakota
University of Southern California
University of Texas
University of Vermont
University of Washington
US CORPORATIONS (8)
Harris, Inc.
Kodak
Lockheed Missles and Space
Microsoft
Motorola Inc.
Robotic Systems Technology, Inc.
Science Systems and Applications, Inc
Southwest Research Institute
US GOVERNMENT RESEARCH LABORATORIES (8)
Air Force Phillips Laboratory
Aeronomy Laboratory, National Oceanic & Atmospheric Administration
Arnold Engineering Development Center, U.S. Air Force
High Altitude Observatory, National Center for Atmospheric Research
Los Alamos National Laboratory
National Optical Astronomy Observatory
Naval Research Laboratory
Space Environment Center, National Oceanic & Atmospheric Administration
INTERNATIONAL ORGANIZATION (.int)
European Space Agency
FOREIGN COUNTRIES (24)
Argentina (ar)
Australia (au)
Austria (at)
Bulgaria (bg)
Canada (ca)
China (cn)
Croatia (hr)
Finland (fi)
France (fr)
Germany (de)
Hungary (hu)
Japan (jp)
India (in)
Italy (it)
The Netherlands (nl)
New Zealand (nz)
Norway (no)
Peru (pu)
Poland (po)
Spain (es)
Russia (ru)
Slovakia (sk)
Switzerland (ch)
United Kingdom (uk)
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha). Currently, PEM has requested GSFC assistance with compiling for
the DEC Alpha since the GSFC executables run on the Alpha and SwRI's compiled
versions do not.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
HALOE continues to provide outstanding scientific measurements in all
science channels. HALOE was powered off at beginning-of-day today as
scheduled. All engineering parameters are within nominal operating region.
Results of gain and bore sight calibration tests conducted during the last
positive and negative beta cycles indicate no change in instrument
performance. HALOE data has been certified through the end of June 2001.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during
the sunlit parts of orbits nearly every day, and measuring nighttime winds
when power is available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
On 16 August 2001 the MPG participated in the UARS
Technical Team Meeting. This series of meetings
is being held to monitor and coordinate activities
associated with the anticipated close down of science
data collection by UARS instruments.
As far as daily operations are concerned, the processing of
instrument Daily Activity Plans continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 04 August 2001 (Orbit 54106, GMT Day 216) through 10 August
2001 (Orbit 54222, GMT Day 222).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 216/0953 54112 AXIS Off
HALOE 216/2333 54120 Power Off
ACRIM 219/2335 54165 Power Off
SUSIM 219/2335 54165 Power Off
SOLSTICE 219/2335 54165 Power Off
HRDI 222/0213 54197 PB (INST) OFF
A Forward-to-Reverse Yaw Around maneuver (number 101) was successfully executed
in Orbit 54176 beginning 220/16:29:57 and completing at 220/17:07:07 on 08
August 2001. The SOLSTICE instrument was in day only and all other instruments
were in their off or standby mode (including HRDI and PEM). Due to the SA
position at 269 degrees, the effective night length of the spacecraft after the
maneuver was approximately 56.4 minutes. Post maneuver analysis of data
indicate that during the Yaw Around, the end-of-night load bus voltage was 23.8
volts and the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 22.9 and 23.0 percent respectively. The next (reverse-to-
forward) YAW is scheduled for 11 September 2001 (DOY 254).
The beta angle ranged between the angles of 19.9 degrees to 5.8 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 46 degrees on 25 August 2001 (DOY 237).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 to 1.3 degrees C. Additional battery performance data for each
battery on 03 August 2001 and 10 August 2001 are provided in Appendix D..
The clock error ranged from Ð10.5 to 11.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 15
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 08
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
Eight (8) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/221 show the observatory data loss to be 303 hours, 00
minutes, 09.248 seconds. This shows a loss of 04 minutes and 20.992 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 04 minutes and 20.992 seconds occurred over two (2) days:
· 2 minutes and 2.000 seconds on Day 219 due to a Late AOS TTR# 14170
· 2 minutes and 18.992 seconds on Day 221 due to a Late AOS TTR# 14170
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS FOT and technical team conducted a series of thruster firings this week in
order to prime the propulsion system should any deboost activities take place.
The A-C attitude thrusters were fired on Tuesday 07 August 2001 over the course
of four passes. All of the tests went very well. The A-C attitude thrusters
were fired for pulse widths of 40,100 and 280 msec. For each pulse width both a
positive and negative torque command was given for all three axes (giving a
total of six commands for each pulse width). The FOT also executed this set of
six commands twice for each pulse width in order to verify that the same
response was seen for multiple data sets.
All responses were nominal with the exception of the negative roll and positive
yaw firings. On these firings a disturbance was seen in two axes instead of
just one. This lead the team to believe the C3 attitude thruster was under
performing. Upon further investigation the C3 thruster appears to be operating
near 50% of its expected performance. Regardless of this finding, UARS still
has more than enough control of all three axes in order to perform an orbit
adjust.
On Wednesday, 08 August 2001 the FOT executed a 4.5 second burn using the B-D
translation thrusters. This burn caused 4 yaw attitude thruster firings to
occur after the termination of the translation burn in order to maintain
attitude error limits. This sequence is exactly what was desired and expected.
From both the realtime plots and data analysis it appears that the B-D
translation thrusters performed exactly like they did the last time they were
used in 1994.
Finally, on 09 August 2001, the B-D translation thrusters were fired again.
This burn was longer in duration than the previous dayÕs and lasted
approximately 13.5 seconds. The primary objective was to verify that the flight
software would properly swap translation with attitude thrusters as needed
during the same ACS cycle. This happened five times and executed exactly as it
was designed.
After the three tests this week the FOT feels confident that the propulsion
system is working properly. Longer burns will be necessary to provide
calibration data to FDF, however the hardware is ready to support any
activities that will be required. The FOT would also like to thank all those
who helped make this happen in such a short period of time.
MEETINGS
The UARS FOT attended a Command Authorization Meeting (CAM) on 06 August 2001
for the upcoming Attitude and Translation Test Burns.
The UARS FOT attended a Go-NoGo Meeting on 07 August 2001 for the Attitude
Thruster Test.
The UARS FOT attended a Script Review on 07 August 2001 for Translation
Thruster Test.
The UARS FOT attended a Go-NoGo Meeting for the Translation Thruster Test on 08
August 2001.
The UARS FOT attended a Go-NoGo Meeting for the Orbit Adjust on 09 August 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 11 September 2001 (DOY 254). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 10 August 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
54112 216/0953 01-488 PEM/AXIS AXIS 1,2 POWER OFF
54116 216/1716 01-487 NBTR NOOPS
54120 216/2333 01-489 HALOE Power OFF
54143 218/1314 01-490 OBC/CLK Clk Adj 0.0 to -20.6 msec/day
54148 218/2019 01-491 NBTR NBTR PBK / RTS-58
54159 219/1314 01-492 PM ATTITUDE THRUSTER BURN
54159 219/1357 01-492 PM THRUSTER BURN CONT
54160 219/1456 01-492 PM THRUSTER BURN CONT
54160 219/1537 01-492 PM THRUSTER BURN COMPLETE
54160 219/1537 01-493 OBC Clk Adj -20.6 to 0.0 msec/day
54165 219/2335 01-495 ACRIM Power OFF
54165 219/2335 01-496 SUSIM Power OFF
54165 219/2335 01-494 SOLSTICE DAY ONLY
54175 220/1537 01-497 MACS BEGIN YAW
54175 220/1537 01-498 MACS GENASCII TABLE 37
54175 220/1537 01-499 MACS GENASCII TABLE 58
54176 220/1734 01-500 ACS/TAM TAM ALIGNMENT
54176 220/1734 01-497 MACS YAW COMPLETE
54178 220/1924 01-501 PM ENGINEERING BURN
54189 221/1511 01-502 PM ENGINEERING BURN
54197 222/0213 01-503 HRDI PB (INST) OFF
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 10 August 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 10 August 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
54143 218/1223 Clk Adj 0.0 to -20.6 msec/day
54160 219/1603 Clk Adj -20.6 to 0.0 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 10 August 2001
03 August 2001 - GMT Day 215 Beta = 54.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 12.0 0.000 0.0
2 +16.8/-78.4 +6.11/+4.51 +29.2/-14.4 24.2 1.030 21.4
3 +22.4/-50.4 +5.15/+3.87 +28.8/-14.4 24.2 1.020 21.1
10 August 2001 - GMT Day 222 Beta = 5.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/+0.0 12.0 0.000 0.0
2 +16.8/-67.2 +5.15/+3.87 +28.8/-12.8 25.4 1.040 20.5
3 +22.4/-56.0 +4.19/+2.92 +28.0/-12.8 25.4 1.030 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
___________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 03 August 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-046 FOT 54108 216/0436 14170 Generic Late Acquisition
01-047 FOT 54116 216/1636 23990 Equipment prob. @ W/S
01-046 FOT 54151 219/0016 14170 Generic Late Acquisition
01-046 FOT 54194 221/2108 14170 Generic Late Acquisition
01-046 FOT 54208 222/2028 14170 Generic Late Acquisition
01-046 FOT 54209 222/2211 14170 Generic Late Acquisition
01-048 FOT 54209 222/2211 N/A Ephemeris Load Rejected
01-046 FOT 54210 222/2351 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-047 FOT 54116 216/1636 23990 Equipment prob. @ W/S
01-048 FOT 54209 222/2211 N/A Ephemeris Load Rejected
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-045 FOT see list August 01 14170 Generic Late Acquisition
SEVNTFW01/32